OpenStudy (oraclethinktank):

What is the domain of x-6/x+6?

OpenStudy (evoker):

Is that (x-6)/(x+6), if so basically the domain is any x that will yield an answer, in essence for this case there is one number that for x you won't get a solution.

OpenStudy (oraclethinktank):

The equation is f(x) = x - 6 over g(x) = x + 6

OpenStudy (evoker):

OK yeah, so basically you can't divide by 0. So what x would cause that to occur.

OpenStudy (oraclethinktank):

6 or -6?

OpenStudy (evoker):

-6, the 6 is ok, you can have 0 in the numerator just not the denominator.

OpenStudy (evoker):

So your domain will be all real numbers except for -6.